us Econ 3.4.31 you buy a computer directly from the manufacturer for S3037 and agree to repay it in 36 equal installments at 1 52% interest per month on the unpaid b our monthly payment is SRound to two decimal places)
t 1.52% interest per month on the unpaid balance, how much are your monthly payments? How much total interest will
Answer #1

Monthly Payment = rP(1+r)N/[(1+r)N-1]

r = 0.0152

N = 36 months

P = 3037

=> Monthly Payment = 3037*( 0.0152)*(1+ 0.0152)36/((1+ 0.0152)36-1) = $110.16

Hence, Monthly Payment = $110.16

Total Amount paid in 36 months = 110.16*36 = $3965.76

Interest paid = Total amount paid - loan amount = 3965.76 - 3037 = $928.76

Hence, Interest Paid = $928.76